Grossmania gave the following additional information on its national economy 


ITEMS

BILLIONS OF DOLLARS

Consumer expenditure

110

Gross fixed capital formation

30

Exports

50

Imports

30

Government Purchases

30

Increasing stock

150

Indirect taxes

455

Factor incomes going out of the country

50

Factor incomes coming into the country

60

Calculate the following for Grossmania

      a) The Total Domestic Expenditure 

                         b) The Total Final Expenditure 

 

 

 

 


1
Expert's answer
2020-08-17T10:05:33-0400

a) final consumption + gross capital formation (investment in a firm, that is, purchase of machinery, equipment, stocks, place of production) + government spending + net exports (export - import; can be both positive and negative) = 110+30+30+(50-30)=190


b) final consumption + gross capital formation + government spending+icreasing stock+ indirect taxes+ net incomes going out of the country =110+30+30+150+455+(50-60)=765